# Circle bundles with PSC over large manifolds

### v3

@@ 1.1. Background and motivation
Let $X$ be a metric space. The macroscopic dimension, denoted by $\dim_{mc}X$, is the minimal $k$ such that $\mathrm{UW}_{k}(X)<\infty$. In other words, $\dim_{mc}X\leq k$ if there is a continuous map $f:X\to Y^{k}$ to a $k$-dimensional simplicial complex $Y$, such that for all $y\in Y$, $\mathrm{diam}f^{-1}(y)<C$ for some uniform constant $C$.
@@
Note that for a compact metric space $X$, $\dim_{mc}(X)=0$. Therefore, when one refers to the macroscopic dimension of a closed manifold $M$, one typically means the macroscopic dimension of the universal cover $\widetilde{M}$, equipped with a metric obtained from lifting some metric on $M$. As any two metrics on a closed manifold are Lipschitz equivalent, this quantity is independent of the chosen metric on $M$.
@@
Let $M$ be a closed connected $n$-dimensional manifold, admitting a PSC metric. Then the macroscopic dimension of the universal cover of $M$ is at most $n-2$, that is,
$\dim_{mc}\widetilde{M}\leq n-2.$
Intuitively, the conjecture says that a manifold admitting a PSC metric, on large scales, looks like a manifold of at least two dimensions less. The primary evidence in this direction is the following simple observation: Given any closed manifold $M$, consider the manifold $M\times S^{2}_{r}$ with the product metric. Here, $S^{2}_{r}$ denotes a sphere of radius $r$. Then for $r$ small enough, the product metric on $M\times S^{2}_{r}$ has positive scalar curvature. In light of the previous remark, we pass to the universal cover of $M\times S^{2}_{r}$, which is $\widetilde{M}\times S^{2}_{r}$. Since the projection map to $\widetilde{M}$ is continuous and each fiber is $S^{2}_{r}$, we have
@@
and therefore the macroscopic dimension $\dim_{mc}(\widetilde{M\times S^{2}_{r}})\leq n-2$. Conjecture 1 asserts that all PSC manifolds must exhibit this behavior. Gromov’s conjecture has been verified in dimension $3$ [GL83, LM23]. In higher dimensions there are several partial results under various assumptions (cf. [BD16, DD22]).
@@
A compact orientable Riemannian manifold $M$ of dimension $n$ is called $\varepsilon^{-1}$-hyperspherical if there exists a $\varepsilon$-contracting map of positive degree from $M$ to the unit $n$-sphere. Further, $M$ is called enlargeable if for every $\varepsilon>0$, there exists a finite cover of $M$ whch is $\varepsilon^{-1}$-hyperspherical and spin.
The basic example of an enlargeable manifold is the unit circle $S^{1}\cong\mathbf{R}/2\pi\mathbf{Z}$. Indeed, the $n$-sheeted cover of $\mathbf{R}/2\pi\mathbf{Z}$ is $\mathbf{R}/2\pi n\mathbf{Z}$. The map $\mathbf{R}/2\pi n\mathbf{Z}\to\mathbf{R}/2\pi\mathbf{Z}$ given as scaling by $1/n$ is a $1/n$-contracting map, for any $n\geq 1$. Also, $S^{1}$ is certainly a spin manifold. It is straightforward to see that
@@
    A manifold $M$ admitting a positive degree map to an enlargeable manifold is also enlargeable, provided $M$ admits a finite spin covering.
@@
In the negative direction we point out the recent comprehensive work due to He [He23] on PSC metrics on non-trivial circle bundles. In particular, He shows that circle bundles with homologically non-trivial fibers do not admit PSC metrics if the base space is a $1$-enlargeable1111111-enlargeable means that the maps to $S^{n}$ in the definition of enlargeability are of degree $1$. manifold of dimension $n\leq 6$ and the circle fibers are homologically non-trivial in the total space [He23, Proposition 4.12].
@@ 1.2. Main results
For $n\geq 2$, in this article, we construct infinitely many examples of $2n$-manifolds $Z$, arising as Donaldson divisors (see, Theorem 1) in symplectic $(2n+2)$-manifolds, such that $Z$ does not admit a PSC metric and macroscopic dimension of $\widetilde{Z}$ is exactly $2n$. Nevertheless, we construct circle bundles over these $Z$,
[…]
such that the $(2n+1)$-dimensional total space $E_{Z}$ admits a PSC metric. An intriguing feature of these examples is that $\dim_{mc}(\widetilde{E_{Z}})\leq n+1$. Thus, even though $E_{Z}$ is obtained from $Z$ by taking a twisted product with $S^{1}$, strangely the macroscopic dimension drops by half for $E_{Z}$. For $n=2$, note that this is sharply in consonance with Gromov’s macroscopic dimension conjecture (Conjecture 1), i.e., $\dim_{mc}(\widetilde{E_{Z}})\leq\dim E_{Z}-2$. We show this using the natural deformation retraction from $\widetilde{E_{Z}}$ to an $(n+1)$-dimensional simplicial complex arising as the spine of the complement of the Donaldson divisor $Z$ in the ambient symplectic $(2n+2)$-manifold.
Note that this phenomenon of the macroscopic dimension decreasing upon taking a (twisted) product with $S^{1}$ is not possible in the case of a trivial product. Indeed,
$\dim_{mc}(\widetilde{Z\times S^{1}})=\dim_{mc}(\widetilde{Z}\times\mathbf{R})=\dim_{mc}(\widetilde{Z})+1.$
@@
[…]
@@
    The macroscopic dimension of the total space $E_{Z}$ is at most $n+1$.
@@
    Let $(M^{2n+2},\omega)$ be a symplectically aspherical [^2] See Definition 4. In particular, any symplectic manifold with $\pi_{2}=0$ (e.g. $T^{2n}$) is symplectically aspherical. manifold with an amenable fundamental group. Then the macroscopic dimension of $Z^{2n}$ is exactly $2n$.
@@
We emphasize that the Conclusions $(1)$ and $(2)$ does not rely on any assumptions either on the fundamental groups of $M$ and $E_{Z}$, or on the (non-)existence of spin or almost spin structures on the universal cover of $E_{Z}$. In particular, Conclusion $(2)$ does not follow from any existing results on Gromov’s conjecture (Conjecture 1).
On the other hand, Conclusions $(3)$ and $(4)$ rest on Proposition 5, where we will show that there is a positive degree map from $Z$ to $M$. Then, Conclusion $(3)$ follows from combining Proposition 7 with a result of Dranishnikov [Dra11], and Conclusion $(4)$ follows from the fact that a manifold admitting a positive degree map to an enlargeable manifold is also enlargeable, provided it admits a finite cover that is spin. In this case, it follows from the classical result of Gromov and Lawson [GL80] that $Z$ does not admit a PSC metric. Finally, it follows from Proposition 4 that varying the parameter $k$ produces an infinite collection of such manifolds $Z$ satisfying the conclusions of Theorem A.
Conclusion $(4)$ of Theorem A gives the examples sought after in Gromov’s question (Question 1) in all even dimension $\geq 4$. The case of odd dimensions $>4$ follows upon considering the bundle $p\times\mathrm{id}:E_{Z}\times S^{1}\to Z\times S^{1}$ obtained by taking product of the bundle $p:E_{Z}\to Z$ with $S^{1}$, and observing that $Z\times S^{1}$ is enlargeable if $Z$ is enlargeable.
@@
Let $M$ be a closed connected $3$-manifold that is either enlargeable or has macroscopic dimension $3$. Let $p:E\to Z$ be a circle bundle. Then the total space $E$ does not admit a metric of positive scalar curvature.
@@ Acknowledgements
The authors would like to thank Shihang He for several helpful comments on the first version of this paper. They would also like to thank Alexey Balitskiy, Alexander Dranishnikov, and Larry Guth for answering their questions and providing several helpful remarks. The authors are also indebted to Bernhard Hanke for his invaluable input on an earlier version of the paper which helped greatly simplify the proofs, as well as for communicating a proof of Theorem B. The second author wishes to thank his advisor Mahan Mj for his interest, as well as his friends and colleagues Ritwik Chakraborty and Sekh Kiran Ajij for several helpful conversations. The second author is supported by the Department of Atomic Energy, Government of India, under project no.12-R&D-TFR-5.01-0500.
@@ 2. Donaldson divisors: definition and properties
Let $(X,\omega)$ be a symplectic manifold. The symplectic form $\omega$ is said to be integral if the cohomology class $[\omega]\in H^{2}(X;\mathbf{R})$ lies in the image of the change of coefficients homomorphism $H^{2}(X;\mathbf{Z})\to H^{2}(X;\mathbf{R})$.
@@
Any symplectic manifold $(X,\eta)$ admits a symplectic form $\omega$ which is integral. To see this, choose a collection of closed $2$-forms on $X$ representing a basis of the vector space $H^{2}(X;\mathbf{R})$, and express the class $\eta$ as an $\mathbf{R}$-linear combination of the chosen closed $2$-forms. Since $H^{2}(X;\mathbf{Q})\subset H^{2}(X;\mathbf{R})$ is dense, we may approximate the real coefficients by rational numbers and then multiply by an integer to clear out the denominators. Let $\omega$ be the resulting closed $2$-form. Then $[\omega]\in H^{2}(X;\mathbf{Z})$. Since non-degeneracy is both an open condition and is scale-invariant, $\omega$ is a non-degenerate $2$-form. Therefore, $\omega$ is an integral symplectic form, as desired.
@@
[Don96, Theorem 1] Suppose $(X^{2n},\omega)$ is a symplectic manifold where $\omega$ is integral. Then, for sufficiently large $k\gg 1$, there exists a symplectic submanifold $Z\subset X$ of codimension $2$, such that $Z$ represents the class Poincaré dual to $k[\omega]$.
@@ 2.1. Complement of Donaldson divisors
[Don96, Proposition 39] The inclusion $j:Z^{2n-2}\hookrightarrow X^{2n}$ induces an isomorphism $j_{*}:\pi_{k}(Z)\to\pi_{k}(X),$ for all $k\leq n-2$, and a surjection for $k=n-1$.
In fact, the following lemma can be deduced from the proof of Proposition 2 in [Don96, pg. 700-701]:
@@
The complement $X^{2n}\setminus Z$ deformation retracts to a simplicial complex of dimension $n$.
@@
By the construction of $Z\subset X$ summarized above, there is an asymptotically holomorphic section $s:X\to L^{\otimes k}$ such that $Z=s^{-1}(\mathbf{0})$. Let $\psi:X\setminus Z\to\mathbf{R}$ be defined by $\psi:=-\log|s|^{2}$. During the course of the proof of [Don96, Proposition 39], it is shown that $\psi$ is a Morse function on $X\setminus Z$ with critical points having index at most $n$. Therefore, by Morse theory, $X\setminus Z$ deformation retracts to an $n$-dimensional simplicial complex. ∎
@@
For Part $(2)$, observe that the Morse function $\psi$ constructed in Lemma 2.2 gives a handlebody decomposition of $W$ by handles of index at most $n$. By reversing the handlebody decomposition, we get that $W$ can be constructed from $\partial W$ by attaching handles of index at least $n$. If $n\geq 3$, then we conclude that all these handles are of index at least $3$. Since $\pi_{1}$ is unchanged under attachements of handles of index bigger than $2$, we conclude $\partial W\hookrightarrow W$ induces an isomorphism on $\pi_{1}$. ∎
@@ 2.2. Topology of Donaldson divisors
In the next proposition, we show that Donaldson divisors of a symplectic manifold $(X^{2n},\omega)$ with integral symplectic form $\omega$, realizing the homology classes Poincaré dual to $k[\omega]$, are topologically distinct for distinct values of $k\gg 1$. We do this by calculating their Euler characteristic.
@@
Let $Z_{k}\subset(X^{2n},\omega)$ denote a Donaldson divisor with $[Z_{k}]=\mathrm{PD}(k[\omega])$. Then, for sufficiently large values of $k\gg 1$, $|\chi(Z_{k})|\sim O(k^{n})$ is monotonically increasing in $k$. In particular, for $k,l\gg 1$ sufficiently large, $Z_{k}\cong Z_{l}$ if and only if $k=l$.
@@
$c_{n-1}(TZ_{k})=\sum_{i=0}^{n-1}(-)^{i}k^{i}\cdot j^{*}c_{n-1-i}(TX)\smile[\omega]^{\smile i},$ (2.4)
where $c_{0}(TX)=1$ by definition. Since the top Chern class is the Euler class, evaluating the cohomology classes appearing in either sides of (2.4) on $[Z_{k}]=k\cdot\mathrm{PD}[\omega]$ we get:
@@
$=\sum_{i=1}^{n}(-1)^{i-1}k^{i}\int_{X}c_{n-1}(TX)\smile[\omega]^{i}$
Note that $[\omega]$ and $c_{i}(TX)$, $0\leq i\leq n$ are all independent of $k$. Hence, $\chi(Z_{k})$ is a polynomial in $k$ with degree $n+1$, with coefficients independent of $k$, and the leading coefficient being nonzero. Therefore, for $k\gg 1$, $|\chi(Z_{k})|$ is monotonically increasing in $k$, as required. ∎
@@
2.3. Donaldson divisors of $M\times S_{g}$
Let $S_{g}$ be a closed surface with genus $g\geq 0$. Let $(M^{2n},\eta_{0})$ be an integral symplectic manifold. We equip the manifold $M\times S_{g}$ with the integral symplectic form $\omega=\eta_{0}\oplus\omega_{0}$, where $\omega_{0}$ denotes a positive, integral area form on $S_{g}$. We now make the following observation, which will be crucial to the proof of Part (4) of Theorem A.

### v5

@@ 1.1. Background and motivation
[Gro96, Dra13] Let $X$ be a metric space. One says $\dim_{\mathrm{mc}}(X)\leq k$ (resp. $\dim_{\mathrm{MC}}(X)\leq k$) if there is a continuous (resp. Lipschitz) map $f:X\to Y^{k}$ to a $k$-dimensional simplicial complex $Y$ equipped with the uniform metric [^1] The uniform metric on a simplicial complex $K$ is one with respect to which all the $n$-simplices of $K$ are isometric to the standard Euclidean $n$-simplex $\Delta^{n}$, for all $n$., such that for all $y\in Y$, $\mathrm{diam}(f^{-1}(y))<C$ for some uniform constant $C$.
@@
Note that it is immediate that $\dim_{\mathrm{mc}}(X)\leq\dim_{\mathrm{MC}}(X)$ [^2] In general, these two notions may not agree even for universal cover of closed manifolds, see [Dra13].. For ease of exposition and simplicity, in this article we shall mostly concern ourselves with the slightly better behaved quantity $\dim_{\mathrm{MC}}(X)$ and refer to it as the macroscopic dimension of $X$, except on occasions where we explicitly state otherwise.
[…]
@@
Let $M$ be a closed connected $n$-dimensional manifold, admitting a PSC metric. Then the macroscopic dimension of the universal cover of $M$ is at most n2n-2, that is,
$\mathrm{dim}_{\mathrm{mc}}(\widetilde{M})\leq n-2.$
Intuitively, the conjecture says that a manifold admitting a PSC metric, on large scales, looks like a manifold of at least two dimensions less. The primary evidence in this direction is the following simple observation: Given any closed manifold $M^{n-2}$, consider the manifold $M^{n-2}\times S^{2}_{r}$ with the product metric. Here, $S^{2}_{r}$ denotes a sphere of radius $r$. Then for $r$ small enough, the product metric on $M\times S^{2}_{r}$ has positive scalar curvature. In light of the previous remark, we pass to the universal cover of $M^{n-2}\times S^{2}_{r}$, which is $\widetilde{M}\times S^{2}_{r}$. Since the projection map to $\widetilde{M}$ is Lipschitz and each fiber is $S^{2}_{r}$, we have
@@
and the macroscopic dimension $\dim_{\mathrm{mc}}(\widetilde{M\times S^{2}_{r}})\leq\dim_{\mathrm{MC}}(\widetilde{M\times S^{2}_{r}})\leq n-2$. Conjecture 1 asserts that all PSC manifolds must exhibit this behavior. Gromov’s conjecture was verified for $3$-manifolds in [GL83]. See also recent works [CL24, LM23] where a stronger version of Gromov’s conjecture is established for $3$-manifolds. In higher dimensions there are several partial results under various assumptions (cf. [BD16, DD22]).
@@
A compact orientable Riemannian manifold $M$ of dimension $n$ is called $\varepsilon^{-1}$-hyperspherical if there exists a $\varepsilon$-contracting map of positive degree from $M$ to the unit $n$-sphere. Further, $M$ is called enlargeable if for every $\varepsilon>0$, there exists a finite cover of $M$ which is $\varepsilon^{-1}$-hyperspherical and spin.
The basic example of an enlargeable manifold is the unit circle $S^{1}\cong\mathbf{R}/2\pi\mathbf{Z}$. Indeed, it is spin and the $n$-sheeted cover of $\mathbf{R}/2\pi\mathbf{Z}$ is $\mathbf{R}/2\pi n\mathbf{Z}$. The map $\mathbf{R}/2\pi n\mathbf{Z}\to\mathbf{R}/2\pi\mathbf{Z}$ given as scaling by $1/n$ is a $1/n$-contracting map, for any $n\geq 1$. It is straightforward to see that
@@
    A manifold $M$ admitting a positive degree Lipschitz map to an enlargeable manifold is also enlargeable, provided $M$ admits a finite spin covering.
@@
In the negative direction we point out the recent comprehensive work due to He [He25] on PSC metrics on non-trivial circle bundles. In particular, He shows that circle bundles with homologically non-trivial fibers do not admit PSC metrics if the base space is a 11-enlargeable33311-enlargeable means that the maps to $S^{n}$ in the definition of enlargeability are of degree $1$. manifold of dimension $n\leq 6$ and the circle fibers are homologically non-trivial in the total space [He25, Proposition 4.12].
@@ 1.2. Main results
For $n\geq 2$, in this article, we construct infinitely many examples of $2n$-manifolds $Z$, arising as Donaldson divisors (see, Theorem 1) in symplectic $(2n+2)$-manifolds, such that $Z$ does not admit a PSC metric and macroscopic dimension of $Z$ is exactly $2n$. Nevertheless, we construct circle bundles over these $Z$,
$S^{1}\hookrightarrow E_{Z}\xrightarrow{p}Z,$
such that the $(2n+1)$-dimensional total space $E_{Z}$ admits a PSC metric. An intriguing feature of these examples is that $\dim_{\mathrm{MC}}(\widetilde{E_{Z}})\leq n+1$. Thus, even though $E_{Z}$ is obtained from $Z$ by taking a twisted product with $S^{1}$, strangely the macroscopic dimension drops by half for $E_{Z}$. For $n=2$, note that this is sharply in consonance with Gromov’s macroscopic dimension conjecture (Conjecture 1), i.e., $\dim_{\mathrm{MC}}(\widetilde{E_{Z}})\leq\dim E_{Z}-2$. We show this using the natural deformation retraction from $\widetilde{E_{Z}}$ to an $(n+1)$-dimensional simplicial complex arising as the spine of the complement of the Donaldson divisor $Z$ in the ambient symplectic $(2n+2)$-manifold.
We remark that even though we state our results with $\dim_{\mathrm{MC}}$, similar techniques can be used to produce $2n$-manifolds $Z$ with $\dim_{\mathrm{mc}}(\widetilde{Z})=2n$, which admit circle bundles whose total spaces exhibit an analogous macroscopic dimension drop phenomenon. We provide such examples in Example 16. Note that this phenomenon of the macroscopic dimension $\dim_{\mathrm{mc}}$ decreasing upon taking a (twisted) product with $S^{1}$ is not possible in the case of a trivial product. Indeed, [Dra13, Theorem 5.7] shows that if $\dim_{\mathrm{mc}}(\widetilde{Z})=2n$, then
$\dim_{\mathrm{mc}}(\widetilde{Z\times S^{1}})=2n+1$
@@
For $n\geq 2$, let $(M^{2n+2},\omega)$ be a closed integral symplectic manifold. Let $Z^{2n}$ be a Donaldson divisor of $M$ with $Z]=\mathrm{PD}(k[\omega])$, for some sufficiently large integer $k\gg 1$. Then there is a circle bundle $S^{1}\hookrightarrow E_{Z}\xrightarrow{p}Z,$ such that the following statements hold:
@@
    The macroscopic dimension of $E_{Z}$ is at most $n+1$, i.e. $\dim_{\mathrm{MC}}(\widetilde{E_{Z}})\leq n+1$.
@@
    Let $(M^{2n+2},\omega)$ be a symplectically aspherical [^4] See Definition 4. In particular, any symplectic manifold with $\pi_{2}=0$ (e.g. $T^{2n}$) is symplectically aspherical. manifold with an amenable fundamental group. Then the macroscopic dimension of $Z^{2n}$ is exactly $2n$, i.e. $\dim_{\mathrm{MC}}(\widetilde{Z})=2n$.
@@
We emphasize that the Conclusions $(1)$ and $(2)$ do not rely on any assumptions either on the fundamental groups of $M$ and $E_{Z}$, or on the (non-)existence of spin or almost spin structures on the universal cover of $E_{Z}$. In particular, Conclusion $(2)$ does not follow from any existing results on Gromov’s conjecture (Conjecture 1).
Conclusion $(3)$ follows from combining Proposition 7 with a result of Dranishnikov [Dra11, Theorem 1.6]. Conclusion $(4)$ follows from Proposition 5 and the fact that a manifold admitting a positive degree Lipschitz map to an enlargeable manifold is also enlargeable, provided it admits a finite cover that is spin. Then, it follows from the classical result of Gromov and Lawson [GL80] that $Z$ does not admit a PSC metric. Finally, it follows from Proposition 4 that varying the parameter $k$ produces an infinite collection of such manifolds $Z$ satisfying the conclusions of Theorem A.
Conclusion $(4)$ of Theorem A gives the examples sought after in Gromov’s question (Question 1) in all even dimension $\geq 4$. The case of odd dimensions >4>4 follows upon considering the bundle $p\times\mathrm{id}:E_{Z}\times S^{1}\to Z\times S^{1}$ obtained by taking product of the bundle $p:E_{Z}\to Z$ with $S^{1}$, and observing that $Z\times S^{1}$ is enlargeable if $Z$ is enlargeable.
@@
Let $M$ be a closed connected $3$-manifold that is either enlargeable or has macroscopic dimension $3$. Let $p:E\to{M}$ be a circle bundle. Then the total space $E$ does not admit a metric of positive scalar curvature.
@@ Acknowledgements
The authors would like to thank Shihang He for several helpful comments on the first version of this paper, as well as for pointing out an error in a previous proof of Theorem B. They would also like to thank Alexey Balitskiy, Alexander Dranishnikov, and Larry Guth for answering their questions and providing several helpful remarks. The authors are also indebted to Bernhard Hanke for his invaluable input on an earlier version of the paper which helped simplify the proofs, as well as for communicating a proof of Theorem B. They also wish to thank the anonymous referees for pointing out inaccuracies as well as several suggestions that have improved this paper. The second author wishes to thank his advisor Mahan Mj for his interest, as well as his friends and colleagues Ritwik Chakraborty and Sekh Kiran Ajij for several helpful conversations. The second author is supported by the Department of Atomic Energy, Government of India, under project no.12-R&D-TFR-5.01-0500.
@@ 2. Donaldson divisors: definition and properties
Let $(X,\omega)$ be a symplectic manifold. The symplectic form $\omega$ is said to be integral if the cohomology class $\omega]\in H^{2}(X;\mathbf{R})$ lies in the image of the change of coefficients homomorphism $H^{2}(X;\mathbf{Z})\to H^{2}(X;\mathbf{R})$.
@@
Any symplectic manifold $(X,\eta)$ admits a symplectic form $\omega$ which is integral. To see this, choose a collection of closed $2$-forms on $X$ representing a basis of the vector space $H^{2}(X;\mathbf{R})$, and express the class $\eta$ as an $\mathbf{R}$-linear combination of the chosen closed $2$-forms. Since $H^{2}(X;\mathbf{Q})\subset H^{2}(X;\mathbf{R})$ is dense, we may approximate the real coefficients by rational numbers and then multiply by an integer to clear out the denominators. Let $\omega$ be the resulting closed $2$-form. Then $\omega]\in H^{2}(X;\mathbf{Z})$. Since non-degeneracy is both an open condition and is scale-invariant, $\omega$ is a non-degenerate $2$-form. Therefore, $\omega$ is an integral symplectic form, as desired.
@@
[Don96, Theorem 1] Suppose $(X^{2n},\omega)$ is a closed symplectic manifold where $\omega$ is integral. Then, for sufficiently large $k\gg 1$, there exists a symplectic submanifold $Z\subset X$ of codimension $2$, such that $Z$ represents the class Poincaré dual to $k[\omega]$.
@@ 2.1. Complement of Donaldson divisors
[Don96, Proposition 39] The inclusion $j:Z^{2n-2}\hookrightarrow X^{2n}$ induces an isomorphism $j_{*}:\pi_{k}(Z)\to\pi_{k}(X),$ for all $k\leq n-2$, and a surjection for k=n1k=n-1.
In fact, the following lemma can be deduced from the proof of Proposition 2 in [Don96, pg. 700-701].
@@
The complement $X^{2n}\setminus Z$ admits a handlebody decomposition consisting of handles of index at most $n$. In particular, it deformation retracts to a simplicial complex of dimension $n$.
@@
By the construction of $Z\subset X$ summarized above, there is an asymptotically holomorphic section $s:X\to L^{\otimes k}$ such that $Z=s^{-1}(\mathbf{0})$. Let $\psi:X\setminus Z\to\mathbf{R}$ be defined by $\psi:=-\log|s|^{2}$. During the course of the proof of [Don96, Proposition 39], it is shown that $\psi$ is a function on $X\setminus Z$ with critical points having index at most $n$. Therefore, by Morse theory, $X\setminus Z$ admits a handlebody decomposition consisting of handles of index at most $n$. Moreover, $X\setminus Z$ deformation retracts to the subcomplex of dimension at most $n$ given by union of the cores of the handles. ∎
@@
For Part $(2)$, recall from Lemma 2.2 that $W$ admits a handlebody decomposition consisting of handles of index at most $n$. By reversing the handlebody decomposition, we get that $W$ can be constructed from $\partial W$ by attaching handles of index at least $n$. If $n\geq 3$, then we conclude that all these handles are of index at least $3$. Since $\pi_{1}$ is unchanged under attachments of handles of index bigger than $2$, we conclude $\partial W\hookrightarrow W$ induces an isomorphism on $\pi_{1}$. ∎
@@ 2.2. Topology of Donaldson divisors
In the next proposition, we show that Donaldson divisors of a closed symplectic manifold $(X^{2n},\omega)$ with integral symplectic form $\omega$, realizing the homology classes Poincaré dual to $k[\omega]$, are topologically distinct for distinct values of $k\gg 1$. We do this by calculating their Euler characteristic.
@@
Let $Z_{k}\subset(X^{2n},\omega)$ denote a Donaldson divisor with $Z_{k}]=\mathrm{PD}(k[\omega])$. Then, for sufficiently large values of $k\gg 1$, $|\chi(Z_{k})|\sim O(k^{n})$ is monotonically increasing in $k$. In particular, for $k,l\gg 1$ sufficiently large, $Z_{k}\cong Z_{l}$ if and only if $k=l$.
@@
$c_{n-1}(TZ_{k})=\sum_{i=0}^{n-1}{(-1)}^{i}k^{i}\cdot j^{*}c_{n-1-i}(TX)\smile[\omega]^{\smile i},$ (2.4)
where $c_{0}(TX)=1$ by definition. Since the top Chern class is the Euler class, evaluating the cohomology classes appearing in either sides of (2.4) on $Z_{k}]=k\cdot\mathrm{PD}[\omega]$ we get:
@@
$=\sum_{i=1}^{n}(-1)^{i-1}k^{i}\int_{X}{c_{n-i}(TX)}\smile[\omega]^{i}$
Note that $\omega]$ and $c_{i}(TX)$, $0\leq i\leq n$ are all independent of $k$. Hence, $\chi(Z_{k})$ is a polynomial in $k$ with degree $n$, with coefficients independent of $k$, and the leading coefficient being nonzero. Therefore, for $k\gg 1$, $|\chi(Z_{k})|$ is monotonically increasing in $k$, as required. ∎
@@
2.3. Donaldson divisors of ${N}\times S_{g}$
Let $S_{g}$ be a closed surface with genus $g\geq 0$. Let $({N}^{2n},\eta_{0})$ be a closed integral symplectic manifold. We equip the manifold ${N}\times S_{g}$ with the integral symplectic form $\omega=\eta_{0}\oplus\omega_{0}$, where $\omega_{0}$ denotes a positive, integral area form on $S_{g}$. We now make the following observation, which will be crucial to the proof of Part (4) of Theorem A.